# Long-Context Generalization with Sparse Attention

### v1

@@ Appendix A $\alpha$-entmax Transformation
$\alpha\text{-entmax}(\bm{z})=\arg\max_{\bm{p}\in\triangle_{n}}\Bigl{(}\bm{p}^{\top}\bm{z}+H_{\alpha}(\bm{p})\Bigr{)},\quad\triangle_{n}=\bigl{\{}\bm{p}\in\mathbb{R}^{n}_{\geq 0}:\sum_{i}p_{i}=1\bigr{\}},$ (12)
where $H_{\alpha}(\bm{p})$ is the Tsallis($\alpha$) entropy. The closed form for $\alpha$-entmax with $\alpha>1$ is
$p_{i}^{\star}=\bigl{[}(\alpha-1)\,z_{i}-\tau(\bm{z})\bigr{]}_{+}^{\tfrac{1}{\alpha-1}},\quad\sum_{i=1}^{n}p_{i}^{\star}=1,\quad[\cdot]_{+}=\max(0,\cdot).$
Here, $\tau(\bm{z})$ is chosen so that $\bm{p}^{\star}$ sums to $1$. Figure 2 illustrates how $\alpha$-entmax with tempered scores $\bm{z}/\theta$ behaves for different choices of $\alpha$.

### v2

@@ Appendix A $\alpha$-entmax Transformation
$\alpha\text{-entmax}(\bm{z}):=\arg\max_{\bm{p}\in\triangle_{n}}\bm{p}^{\top}\bm{z}+H_{\alpha}(\bm{p}),\quad\triangle_{n}:=\{\bm{p}\in\mathbb{R}^{n}\,:\,\bm{p}\geq\mathbf{0},\mathbf{1}^{\top}\bm{p}=1\},$ (12)
where $H_{\alpha}(\bm{p})$ is the Tsallis($\alpha$) entropy (Tsallis, 1988). The closed form for $\alpha$-entmax with $\alpha>1$ is
$p_{i}^{\star}=\bigl{[}(\alpha-1)\,z_{i}-\tau(\bm{z})\bigr{]}_{+}^{\tfrac{1}{\alpha-1}},$
where $[\cdot]_{+}=\max(0,\cdot)$, and $\tau(\bm{z})$ is chosen so that $\bm{p}^{\star}$ sums to $1$. Figure 2 illustrates how $\alpha$-entmax with tempered scores $\bm{z}/\theta$ behaves for different choices of $\alpha$.